OpenStudy (anonymous):

How do I start this problem? d/dx [e^-x ln(x + 1)]

OpenStudy (johnweldon1993):

Well you have 2 functions being multiplied and you want the derivative...sounds like a product rule here

OpenStudy (johnweldon1993):

\[\large \frac{d}{dx}(u,v) = \frac{du}{dx}v + \frac{dv}{dx}u\] Letting u = e^-x and v = ln(x + 1)

OpenStudy (anonymous):

ok. thank you